OpenStudy (anonymous):

What is the y-intercept of the line passing through the point (9, -7) with a slope of -1/8 ? A 47/8 B -47/8 c 65/8 D -65/8

OpenStudy (ganpat):

Equation of line is y =mx + C Substitute value of (x,y) as (9,-7) and slope as m = (-1/8) to find out C for Y intercept, put x=0, and put values of m = (-1/8) and C

OpenStudy (ganpat):

y = mx +c (x,y) = (9,-7) and m = -(1/8) -7 = -9/8 + C C = 9/8 + 7 c = 65/8

OpenStudy (ganpat):

So now you have m = -(1/8) C = 65/8 and x = 0... to find Y intercept using the same equation y = mx + c